The perimeter of a rectangle is 72 inches. If the length of the rectangle is nine inches more than two times the width, find the dimensions.
Answer by JulietG(1812)

Perimeter (fence around) = 2L + 2W
L = 2W + 9
Substitute the value of L from the bottom equation into the top.
72 = 2L + 2W
72 = 2(2W+9) + 2W
Distribute
72 = 4W + 18 + 2W
Add
72 = 6W + 18
Subtract 18 from each side
54 = 6W
Divide each side by 6
9 = W
.
If the width is 9, then the length is (2*9)+9 = 27
27 + 27 + 9 + 9 = 72 (perimeter = 2*length + 2*width)